OpenStudy (firejay5):

The initial # of bacteria in a culture is 12,000. The # after 4 hours of the application of a new drug is 4,000. 1. Write an exponential function to model the population y of bacteria after x hours. 2. How many bacteria are there after 24 hours?

OpenStudy (firejay5):

How would I change this: to exponential function of y = A * B^x 1. y=12000e^(kt) 2. 4000=12000e^4k e^4k=4000/12000=1/3 4k=ln(1/3) therefore k=[ln(1/3)]/4 To find how many bacteria after 24hours. p=12000e^{24x[ln(1/3)]/4} and you'll get 16.46 population

OpenStudy (firejay5):

@Hero @ikram002p @TheSmartOne @aaronq @undeadknight26